OpenStudy (anonymous):

You are given the following function. f(t) = t + √t Find the derivative of the function using the definition of derivative.

terenzreignz (terenzreignz):

Yikes... \[\Large f(t) = t + \sqrt t\] \[\Large f'(t) = \lim_{h\rightarrow 0}\frac{f(t+h)-f(t)}{h}\] This will not be pretty lol

terenzreignz (terenzreignz):

I'm sure you've gotten at least that^ far. What you have got now? ^_^

OpenStudy (anonymous):

\[1 + \lim_{h \rightarrow 0} \frac{ \sqrt{t + h} - \sqrt{t}}{ h } \] what's next?

terenzreignz (terenzreignz):

Hmmnn...

terenzreignz (terenzreignz):

argh... this is why nobody does it this way XD... hang on

terenzreignz (terenzreignz):

Okay, try multiplying with this: \[\Large \frac{\sqrt{t+h}+\sqrt t}{\sqrt {t+h}+ \sqrt t}\]

terenzreignz (terenzreignz):

@swedishlad bet you already got there before me xD

OpenStudy (anonymous):

\[1 +\lim_{h \rightarrow 0}\frac{ h }{ h(\sqrt{t+h}+\sqrt{t}) }\]

terenzreignz (terenzreignz):

You can cancel out the h now \[\Large 1 +\lim_{h \rightarrow 0}\frac{ \cancel h }{ \cancel h(\sqrt{t+h}+\sqrt{t}) }\]

terenzreignz (terenzreignz):

And you'd be done lol

OpenStudy (anonymous):

\[\frac{ 1 }{ 2\sqrt{t}} + 1\]

OpenStudy (anonymous):

boom, thanks man

terenzreignz (terenzreignz):

No problem ^_^